OpenStudy (anonymous):

A 0.053-kg arrow is fired horizontally. The bowstring exerts an average force of 70 N on the arrow over a distance of 0.90 m. With what speed does the arrow leave the bow? A 0.053-kg arrow is fired horizontally. The bowstring exerts an average force of 70 N on the arrow over a distance of 0.90 m. With what speed does the arrow leave the bow? @Physics

OpenStudy (anonymous):

is it greater than 1000 m/s?

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

\[\Large \begin{array}{l} W = Fd\cos \theta \\ = 0.9 \times 70\cos 0 = 6.3J\\ KE = \frac{1}{2}m{v^2}\\ {v^2} = \frac{{2KE}}{m}\\ v = \sqrt {\frac{{2KE}}{m}} \\ = \sqrt {\frac{{2 \cdot 6.3}}{{0.053}}} \\ \cong 15.42m{s^{ - 1}} \end{array}\]